# Deep Learning-based MRI Reconstruction with Artificial Fourier Transform Network (AFTNet)

### v2

@@ 3.4 Experimental data
Two datasets were used in this study: a complex-valued normal-field human brain MRI dataset from the fastMRI dataset[22] and a complex-valued human brain MRS dataset from the Big GABA dataset[23]. The proposed methods were trained on these datasets separately.
The normal-field human brain MRI dataset contains fully sampled brain MRIs obtained on 3 and 1.5 Tesla magnets. We selected 4-channels axial T1-weighted and T2-weighted scans from the raw fastMRI dataset. A total number of 943 scans were used with 794, 99, and 50 each for the training, validation, and test set. All the scans were first normalized to the max intensity value of one and cropped to $640\times 320$ matrix size.
The human brain MRS dataset contains GABA-edited MEGA-PRESS data obtained on 3T Philips scanners from different sites. Each subject was scanned for 320 averages (160 ON and 160 OFF repetitions). The data points acquired by each repetition is 2048. A total number of 101 scans were selected from the Big GABA dataset with 80, 10, and 11 each for the training, validation, and test set. All the scans were first normalized to the max spectra magnitude value of one.
During the accelerated MRI reconstruction, all the k-space data was undersampled from the fully sampled k-space by applying a mask in the phase-encoding direction. We use the acceleration rate (or acceleration factor), to denote the level of scan time reduced for the undersampled k-space data, which is defined as the ratio of the amount of k-space data required for a fully sampled image to the amount collected in an undersampled k-space data[32]. The sampling ratio, SR, is also used to denote the information retained in the undersampled k-space data, which is defined as the inverse of the acceleration rate. An equispaced mask with approximate acceleration matching is used to undersample the k-space data. The fraction of low-frequency columns to be retained for acceleration rates 2x, 4x, and 8x are 16%, 8%, and 4% respectively.
During the denoised MRS reconstruction, we use the reduction rate, R, to denote the noise level, which is defined as the ratio of the number of total repetitions (160 for this study) to the number of repetitions retained for a noisy input. We generate the noisy FIDs with 5 reduction rates of 10, 20, 40, 80, and 160.
@@ 3.5 Measurement of Reconstruction Quality
Three metrics were adopted for the quantitative evaluation of the image quality compared with the ground truth: structural similarity (SSIM)[33], peak signal-to-noise ratio (PSNR), and normalized root mean squared error (NRMSE). For the quality measurement of the 1D spectra, another three metrics were used: Pearson correlation coefficient (PCC), Spearman’s rank correlation coefficient (SCC), and goodness-fitting coefficient (GFC)[34]. The GFC is introduced to evaluate the goodness of the mathematical reconstruction with a value ranging from 0 to 1, where 1 indicates a perfect reconstruction. If $\hat{y}_{i}$ is the predicted value of the $i$-th sample and $y_{i}$ is the corresponding true value, then the GFC estimated over $n_{\text{samples}}$ is defined as
@@ 4.1 Comparison study
The effectiveness of the AFTNet was evaluated on the test set of the human normal-field MRI dataset through a comparative study as shown in Figure 4, where the up row shows the visualization of the baseline models, including conventional Fourier transform-based networks and score-MRI. The bottom row shows the proposed networks and ground truth derived from the fully-sampled k-space data. Both the residual maps against the ground truth and the zoomed-in areas are presented. The residual maps are normalized to the max value of $0.5$ such that they are comparable across all models. We first compare the AFT-only network with the results after DFT. It clearly shows that the light trainable AFT block not only approximates the DFT but also facilitates artifact removal with the presence of signal non-ideality. Comparing the results of the AFT-only network and score-MRI, which solves the image reconstruction inverse problem based on score-based generative models, we demonstrate that even with such a simple structure, AFT can achieve similar performance in terms of PSNR and NRMSE. The score-MRI was originally trained on the knee dataset and we fine-tuned it on our dataset. Qualitatively, AFTNet outperformed score-MRI, as can be seen from the difference magnitude map and zoomed-in area in Figure 4 by comparing the results of AFTNet and score-MRI. The quantitative metrics demonstrate the superior performance of AFTNet over score-MRI, as shown in Table 1, including the SSIM, PSNR, and NRMSE across all acceleration rates. The statistical t-test between the metrics of score-MRI and AFTNet also shows the superior accelerated reconstruction of the proposed method with a p-value under 0.0001. Furthermore, the AFT block can serve as a better replacement for DFT by comparing the results of DFTNet and AFTNet, where all AFTNet shows better performance than the DFTNet, as can be seen from the residual maps. It should be pointed out that AFTNet-KI significantly outperforms DFTNet-KI, indicating that AFT can be implemented to connect two networks while preserving the capability of both networks.
Figure 4: Comparison between baseline methods and AFTNet. Here we present the qualitative results under 4x acceleration rate, where a 1D 4x equal-spaced sampling mask with 8% of low-frequency columns retained is implemented to undersample the input k-space data. Residual map shows the difference magnitude against ground truth (in Hot colormap). Yellow and blue boxes show the zoomed-in version of the indicated area. White numbers in the magnitude images indicate PSNR (dB) and SSIM. White numbers in the residual maps indicate NRMSE.
@@ 4.2 Ablation study
Different structures of AFTNet, as mentioned in Section 3.2, were compared under datasets with different field strengths and different modalities to verify the stability and generality of the AFTNet. In addition, the effectiveness of the front-end/back-end convolutional networks is also evaluated in this section. To validate the robustness of AFTNet to k-space artifacts, these proposed AFTNet structures were compared on the image reconstruction and accelerated reconstruction as described in Section 3.4. Furthermore, the extended AFTNet was compared with numerical methods using 1-dimensional MRS FID data on the denoised reconstruction.
Figure 5: Human 1.5/3T MRI reconstruction results. (a) Ground truth, (b) proposed method, (c) difference magnitude of (a) and (b) (in Hot colormap).
First, we show the results of human 1.5/3T MRI reconstruction using raw fully-sampled fastMRI k-space data in Figure 5. All the images shown here and in the following sections are cropped so that the anti-aliasing placed outside the field of view (FOV) in phase-encoding directions is removed. The ground truth image is derived by applying conventional Fourier transformation to the k-space data. It can be seen that the ground truth image obtained from FT is identical to the AFT prediction, which human observers can not distinguish. The results adhere to the mathematical description we discussed in Section 2.1. The residual map (pixel-wise difference between the ground truth image and the AFT prediction) shows that no brain structural information is presented. The grid-like remaining error is mainly caused by precision loss during floating-point calculation in matrix multiplication.
@@
In Figure 4, we show the results of human 1.5/3T accelerated reconstruction using under-sampled fastMRI k-space data. In the first row, we see the reconstructions from 1D 4x equal-spaced sampling, in which 8% of low-frequency columns are retained. Here, we compare different AFTNet structures with the zero-filling method. AFTNet-KI performs outstanding reconstruction, where less structural difference can be seen from the residual map in the second row. The third row shows zoomed-in areas of both images and residual maps. AFTNet-I produces a more blurry reconstruction which loses the structural details. Reconstruction through AFTNet-K induces foggy artifacts, which is reflected in terms of SSIM. Figure 6 shows the accelerated reconstruction results by comparing AFT and AFTNet (I, K, and KI) to their counterparts in terms of SSIM across acceleration rates 2x, 4x, and 8x. The performance of DFT drops linearly as the acceleration rate increases while the AFTNet methods are more robust to the acceleration scale. The t-test between AFT-based and DFT-based structures indicates that the AFTNet overcomes all DFTNet structures significantly except for the I model in the 2x acceleration rate. The results of AFTNet on different acquisition types and system field strength in Figure 7 demonstrate that higher system field strength yields better reconstruction which is due to the original imaging quality between different fields and the reconstruction on T2w images performs better than that on T1w images.
A comprehensive comparison of quantitative metrics on the test set is provided in Table 1 and Table 2 for accelerated reconstruction. AFTNet-KI outperforms other AFTNet structures on all the different acceleration rates and both AFTNet-KI and AFTNet-I perform significantly better than other AFTNet-K structures. Table 2 shows detailed quantitative metrics of human 1.5/3T MRI accelerated reconstruction results grouped by image contrast and system field strength. It is worth mentioning that although AFTNet-K does not outperform other AFTNet structures in the accelerated reconstruction task, it demonstrates the ability to learn in a sparse frequency domain and its sparse representations with a complex-valued convolutional network.
4.3 Human normal-field MRS study
@@
Magnetic resonance spectroscopy, namely MRS, is widely used for measuring human metabolism. While MRS has the potential to be highly valuable in clinical practice, it poses several challenges such as low signal-to-noise ratio, overlapping metabolite signals, experimental artifacts, and long acquisition times. Here, the AFTNet is leveraged as a unified MRS reconstruction approach, which aims to reconstruct and process the FID in parallel, as shown in Figure 8.
We trained our model on the MEGA-PRESS spectra from the Big GABA dataset for two reasons. First, as a proof of concept study, to guarantee the convergence of the supervised learning task, we need the dataset to be sufficient in the number of samples, good in data quality, and publicly available. Thus the Big GABA dataset perfectly meets our requirements. Second, the smaller targeted signals are revealed by the subtraction of 2 spectra containing strong signals (OFF and ON), which provide a good way to verify the performance of the proposed method by measuring the subtraction artifacts. A total number of 101 subjects acquired by the Philips scanners were used in the training. For each subject, a standard GABA ON/OFF edited MRS acquisition was run, where ON editing pulses were placed at 1.9 ppm and OFF editing pulses were placed at 7.46 ppm. The acquisition number is 320 (160 ON and 160 OFF transients) per subject. The AFTNet was trained with an input size of 2048. The ground truth of the ON/OFF spectra is derived by taking the average over 160 acquisitions. We denote the ground truth as noiseless signals. For the training, we combined randomly sampled acquisitions of each subject to retrieve a noisy signal. By decreasing/increasing the number of sampled acquisitions, we can generate signals with higher/lower noise. We use the reduction rate (R) to denote the level of noise, which is defined as the ratio of the total acquisition number and the number of acquisitions sampled. This quantity is very handy to assess the power of denoising methods in practical terms. Retrieving accurate denoised signals at a high R has implications for the potential reduction of total experimental time.
The results of the AFTNet approach and conventional numerical methods with Gaussian line broadening are illustrated in Figure 9. The first row shows the reconstructed spectrum from the numerical methods and the proposed AFTNet. The second row indicates the reconstructed spectrum overlaid with the ground truth. The third row plots the difference between the reconstructed spectrum and the ground truth. Under a reduction rate of 80, where only 2 acquisitions were used over all 160 acquisitions, the AFTNet shows excellent performance at high reduction rates. The AFTNet outperforms other methods for the DIFF spectra, indicating that the AFTNet removes the noise in the FIDs while preserving the subject-level features. We used the Goodness-of-Fit Coefficient (GFC) to measure the similarity between the reconstructed spectra and the ground truth, as shown in the Table 3. The metric value increases as the reduction rate decreases, but the absolute difference between high and low reduction rates is tiny (0.9798 for OFF spectra under a reduction rate of 10 vs. 0.9688 for OFF spectra under a reduction rate of 160). In addition, AFTNet outperforms the DFT+GLB (Gaussian Line Broadening) method across all metrics in the table.
@@ 5 Discussion
In this study, a unified MR image reconstruction framework is proposed, which is composed of two main components: artificial Fourier transform block and complex-valued residual attention U-Net. The AFT block is used to approximate the conventional DFT, which is demonstrated in Table 1 and proofed in Section 3.1. The front-end/back-end convolutional layers are used to extract features at different levels in the k-space/image domains and play different roles in various tasks. As shown in Table 1, both front-end and back-end convolutional layers show superior accelerated reconstruction performance under all sampling ratios compared with single front-end/back-end convolutional layers. This is potentially because the undersampling is performed in k-space where the artifacts are separated from the non-artifact. While in the image domain, it is converted to aliasing overlapped over the whole image. The artifacts removal task can be recast as an image inpainting problem in the k-space domain which can be done more easily by the front-end convolutional layers. However, all the structures with front-end-only convolutional layers show lower performance, indicating that the sparse representation of k-space data makes it harder for a convolutional network to extract noise information in the low-frequency areas and back-end convolutional layers are necessary to achieve the optimal performance.
@@
Domain-transform manifold learning has been introduced for years and several deep learning frameworks were developed based on this idea. The first model, AUTOMAP[11], proposed the simple FC-Conv structure which can only be applied to images with small matrix size due to its redundant FC layers. DOTA-MRI[13] extended AUTOMAP to Conv-FC-Conv structure and applied FC layers to only one dimension (phase-encoding direction). However, it did not extend the model structure and implement end-to-end complex-valued neural networks. The AFTNet we proposed in this study solves the problem mentioned above through a modular-designed AFT block. We also demonstrated that the extended AFTNet can also be applied to 1D data in Section 4.3. In addition, previous works define the loss in the magnitude image, while we calculate the loss in the complex-valued image domain, which preserves the relations between the real and imaginary parts. The phase is then derived from the output of AFTNet, which is essential for several phase-based applications, such as flow quantification and fat–water separation.
@@
Complex-valued neural networks, especially complex-valued convolutional networks[21, 20], have been studied for MRI reconstruction but they mainly focused on simple tasks or only applied it to the image domain. We investigate the different impacts of complex-valued convolutional networks on the k-space and image domain and extend the application to accelerated reconstruction and denoised reconstruction, which are more clinically important. We also incorporate domain-manifold learning by adding domain transform blocks which determine the mapping between the k-space and image domain instead of conventional discrete Fourier transform. It is more robust to noise and signal nonideality due to imperfect acquisition. We also extend the application of complex-valued convolutional networks to 1D MRS denoised reconstruction, which has not been studied in previous work.
@@
One remaining methodological limitation is that the FC layers used by AFTNet narrow the application to datasets with various image matrix sizes. Although the convolutional layers are not sensitive to the image matrix sizes and cropping/padding can be applied to match the desired sizes, the features of FC layers need to be selected carefully which requires further investigation. Another parameter that needs to be taken into account is the coil number. In this study, we selected especially 4-channel MRI data for convenience of data preprocessing. While deep learning-based coil combinations could be incorporated into the framework in future work. Furthermore, diffusion models are shown to be a powerful tool for image reconstruction across body regions and coil numbers[14]. However, the score-MRI we compared in this study does not demonstrate superior performance compared with AFTNet and the inference stage time is extremely long. This is potential because the backbone of the score-MRI is still a real-valued U-Net like structure and the relation between the real and imaginary part is not considered during the calculation of the score function. For future works, the AFTNet could be further extended by leveraging diffusion-based models with complex-valued convolutional networks as the backbone and careful optimization to reduce the inference time.

### v3

@@ 4.2 Experimental data
In this study, two datasets were used: fastMRI dataset [29]) and Big GABA dataset [30]. The proposed methods were trained on these datasets separately.
The fastMRI dataset contains fully sampled brain MRIs obtained on 3 and 1.5 Tesla magnets. We selected 4-channel axial T1 and T2-weighted scans from the raw fastMRI dataset. We applied a data filtering strategy by counting the number of images with similar shapes and field strength, and selecting the groups that have more than 100 samples. This process resulted in 671 T2-weighted images at 1.5T, 157 T2-weighted images at 3T, and 115 T1-weighted images at 3T. All selected images are acquired on Siemens scanner. A total of 943 scans were used with 794, 99, and 50 each for the training, validation, and test set. The detailed scan parameters are listed in Table 1. Data preprocessing includes normalization to the maximum intensity value of one in the image domain and cropping to $640\times 320$. For experiment on accelerated reconstruction, the fully sampled k-space data was undersampled by applying a mask in the phase-encoding direction. We use the acceleration rate (or acceleration factor) to denote the level of scan time reduced for the undersampled k-space data, which is defined as the ratio of the amount of k-space data required for a fully sampled image to the amount collected in an undersampled k-space data [31]. The sampling ratio, SR, is also used to denote the information retained in the undersampled k-space data, which is defined as the inverse of the acceleration rate. An equispaced mask with approximate acceleration matching is used to undersample the k-space data. The fraction of low-frequency columns to be retained for acceleration rates 2x, 4x, and 8x is 16%, 8%, and 4%, respectively.
The in vivo MRS dataset used in this study is from the BIG GABA ([30]) repository and contains GABA-edited MEGA-PRESS data acquired on 3T Philips scanners from multiple sites. The voxel was placed in the medial parietal lobe of healthy subjects aged 18 to 36 years. The data consist of 101 pairs of edit-ON/-OFF spectra, each with 160 ON averages and OFF averages. The AFTNet was trained with an input size of 2048. The ground truth of the ON/OFF spectra is derived by taking the average over 160 transients. We denote the ground truth as noiseless signals. For training, we randomly sampled transients from each subject to reconstruct the averaged signal. By varying the number of transients in computing the average sampled acquisitions, we can generate signals with different levels of signal-to-noise ratio (SNR). We use the term reduction rate (R) to refer to the ratio of the total number of transients and the number of transients used in calculating the average. Being able to reconstruct accurate denoised signals at high R has implications for reducing acquisition time. From a total number of 101 scans in the dataset, 80, 10, and 11 scans were assigned for the training, validation, and test set, respectively. All spectra were normalized to the maximum value of its magnitude. In denoised MRS reconstruction, we use the reduction rate, R, to denote the noise level, which is defined as the ratio of the number of total repetitions (160 for this study) to the number of repetitions retained for a noisy input. We generate noisy free induction decay (FIDs) with 5 reduction rates of 10, 20, 40, 80, and 160.
@@ 4.3 Measurement of Reconstruction Quality
Three metrics were adopted for the quantitative evaluation of image quality compared to ground truth: structural similarity (SSIM) [32], peak signal-to-noise ratio (PSNR), and normalized root mean squared error (NRMSE). For the quality measurement of the 1D spectra, three other metrics were used: Pearson’s correlation coefficient (PCC), Spearman’s rank correlation coefficient (SCC) and goodness fitting coefficient (GFC) [33]. The GFC is introduced to evaluate the goodness of the mathematical reconstruction with a value ranging from 0 to 1, where 1 indicates a perfect reconstruction. If $\hat{y}_{i}$ is the predicted value of the $i$-th sample and $y_{i}$ is the corresponding true value, then the GFC estimated over $n_{\text{samples}}$ is defined as
@@
4.4 Implementation details
We construct a batch size of 1 and optimize the network using the ADAM [34] optimizer. We use a cosine annealing learning rate scheduler [35]. The initial learning rate is $10^{-3}$ and the final learning rate is $10^{-5}$. We set the negative slope of LeakReLU to $0.1$. All methods are trained for 50 epochs. All experiments were performed with PyTorch 2.2.2 and a Quadro RTX 6000 GPU.
For experiment on the MRI accelerated reconstruction, we compared AFTNet with zero-filling (DFT), numerical method GRAPPA [36], real-valued model KIKI-net [37] and diffusion model score-MRI [13]. For experiment on the MRS denoised reconstruction, we compared AFTNet with Gaussian line broadening (GLB).
@@ 5.1 Comparison study
Figure 4: Comparison between baseline methods and AFTNet. Here we present the qualitative results f a T2w image on 3T
under 4x acceleration rate, where a 1D 4x equal-spaced sampling mask with 8% of low-frequency columns retained is implemented to undersample the input k-space data. Residual map shows the difference magnitude against ground truth (in Hot colormap). Yellow and blue boxes show the zoomed-in version of the indicated area. White numbers in the magnitude images indicate PSNR (dB) and SSIM. White numbers in the residual maps indicate NRMSE.
The effectiveness of AFTNet was evaluated on the test set of the human normal-field MRI dataset through a comparative study as shown in Figure 4, where the up row shows the visualization of the baseline models, including conventional Fourier transform-based networks and score-MRI. The bottom row shows the proposed networks and ground truth derived from the fully sampled k-space data. Both the residual maps against the ground truth and the zoomed-in areas are presented. The residual maps are normalized to the maximum value of $0.5$ such that they are comparable across all models. We first compare the AFT-only network with the results after DFT. It clearly shows that the light-trainable AFT block not only approximates the DFT but also facilitates artifact removal with the presence of signal non-ideality. Comparing the results of the AFT-only network and score-MRI, which solves the image reconstruction inverse problem based on score-based generative models, we demonstrate that even with such a simple structure, AFT can achieve similar performance in terms of PSNR and NRMSE. The score-MRI was originally trained on the knee dataset and we fine-tuned it on our dataset. Qualitatively, AFTNet outperformed score-MRI, as can be seen from the difference magnitude map and the zoomed-in area in Figure 4 by comparing the results of AFTNet and score-MRI. Quantitative metrics demonstrate the superior performance of AFTNet over score-MRI and KIKI-net, as shown in Table 1, including SSIM, PSNR, and NRMSE in all acceleration rates. The statistical t-test between the score-MRI and AFTNet metrics also shows the superior accelerated reconstruction of the proposed method with a p-value below 0.0001. Furthermore, the AFT block can serve as a better replacement for DFT by comparing the results of DFTNet and AFTNet, where all AFTNet shows better performance than DFTNet, as can be seen from the residual maps. It should be pointed out that AFTNet-KI significantly outperforms DFTNet-KI, indicating that AFT can be implemented to connect two networks while preserving the capability of both networks.
@@ 5.2 Ablation study
Different AFTNet structures, as mentioned in Section 3.2, were compared in datasets with different field strengths and different modalities to verify the stability and generality of the AFTNet. In addition, the effectiveness of the front-end/back-end convolutional networks is also evaluated in this section. To validate the robustness of AFTNet to k-space artifacts, these proposed AFTNet structures were compared on image reconstruction and accelerated reconstruction as described in Section 4.2. Furthermore, the extended AFTNet was compared with numerical methods using 1-dimensional MRS FID data on the denoised reconstruction.
Figure 5: Fully sampled MRI reconstruction results. Here we present the qualitative results of a T2w image on 1.5T. (a) Ground truth, (b) proposed method, (c) difference magnitude of (a) and (b) (in Hot colormap).
First, we show the results of human 1.5/3T MRI reconstruction using fully sampled fastMRI k-space data in Figure 5. All the images shown here and in the following sections are cropped so that the anti-aliasing placed outside the field of view (FOV) in phase-encoding directions is removed. The ground truth image is derived by applying the conventional Fourier transformation to the k-space data. It can be seen that the ground truth image obtained from FT is identical to the AFT prediction, which human observers can not distinguish. The results adhere to the mathematical description we discussed in Section 2.1. The residual map (pixel-wise difference between the ground truth image and the AFT prediction) shows that no brain structural information is presented. The grid-like remaining error is mainly caused by precision loss during floating-point calculation in matrix multiplication.
@@
In Figure 4, we show the results of MRI accelerated reconstruction using under-sampled fastMRI k-space data. In the first row, we see the reconstructions from the 1D 4x equal space sampling, in which 8% of low-frequency columns are retained (see Figure 1). Here, we compare different AFTNet structures with the zero-filling method. AFTNet-KI performs outstanding reconstruction, where less structural difference can be seen from the residual map in the second row. The third row shows the zoomed-in areas of both the images and residual maps. AFTNet-I produces a more blurry reconstruction which loses the structural details. Reconstruction through AFTNet-K induces foggy artifacts, which are reflected in terms of SSIM. Figure 6 shows the results of the accelerated reconstruction comparing AFT and AFTNet (I, K, and KI) with their counterparts in terms of SSIM in acceleration rates of 2x, 4x and 8x. The performance of DFT drops linearly as the acceleration rate increases, while the AFTNet methods are more robust to the acceleration scale. The t-test between AFT-based and DFT-based structures indicates that AFTNet significantly outperforms all DFTNet structures except the I model in the 2x acceleration rate. The results of AFTNet on different acquisition types and system field strength are shown in Figure 7. The difference of performance is mainly caused by the image quality between the 3T and 1.5T images and the inrisint contrast difference between T2w and T1w images.
A comprehensive comparison of quantitative metrics on the test set is provided in Table 1 and Table 3 for accelerated reconstruction. In addition to the metrics evaluated on the magnitude images, we provide the evaluation on the phase images in Table 2. AFTNet-KI outperforms other AFTNet structures on all the different acceleration rates, and both AFTNet-KI and AFTNet-I perform significantly better than other AFTNet-K structures. Table 3 shows detailed quantitative metrics of the human 1.5/3T MRI accelerated reconstruction results grouped by image contrast and system field strength. It is worth mentioning that although AFTNet-K does not outperform other AFTNet structures in the accelerated reconstruction task, it demonstrates the ability to learn in a sparse frequency domain and its sparse representations with a complex-valued convolutional network.
5.3 Generality of AFTNet
@@
The results of the proposed AFTNet approach and traditional DFT method with and without GLB are illustrated in Figure 8. The first row shows the reconstructed spectra from the methods being compared. The second row shows the reconstructed spectra overlaid with the ground truth. The third row shows the difference between the reconstructed spectra and the ground truth. The reconstructed spectra illustrated here represent a reduction rate of 80, where only two out of a total of 160 transients were used to construct the spectrum. The AFTNet approach outperforms the other methods in all reduction rates listed in the table. We used the GFC to measure the similarity between the reconstructed spectra and the ground truth, as shown in Table 4. Although the GFC decreases as the reduction rate increases, the AFTNet method is able to maintain a high performance even at very high reduction rates. For example, the difference in GFC between the OFF spectra under a reduction rate of 10 and 160 is very small, around 0.9798 and 0.9688, respectively.
@@ 6 Discussion
In this study, a unified MR image reconstruction framework is proposed, which is composed of two main components: artificial Fourier transform block and complex-valued residual attention UNet. The AFT block is used to approximate conventional DFT, which is demonstrated in Table 1 and shown in Section 3.1. The front-end/back-end convolutional layers are used to extract features at different levels in the k-space/image domains and play different roles in various tasks. As shown in Table 1, both front- and back-end convolutional layers show superior accelerated reconstruction performance under all sampling ratios compared to single front-end/back-end convolutional layers. This is potentially because the undersampling is performed in k-space where the artifacts are separated from the non-artifact. While in the image domain, it is converted to an aliasing overlapped over the whole image. The artifact removal task can be recast as an image inpainting problem in the k-space domain which can be done more easily by the front-end convolutional layers. However, all structures with front-end-only convolutional layers show lower performance, indicating that the sparse representation of k-space data makes it harder for a convolutional network to extract noise information in the low-frequency areas and back-end convolutional layers are necessary to achieve the optimal performance.
@@
Domain-transform manifold learning has been introduced for years, and several deep learning frameworks were developed based on this idea. The first model, AUTOMAP [10], proposed the simple FC-Conv structure which can only be applied to images with a small matrix size due to its redundant FC layers. DOTA-MRI [12] extended AUTOMAP to Conv-FC-Conv structure. However, it did not extend the model structure and implement CVNNs. The AFTNet we proposed in this study solves the problem mentioned above through a modular-designed AFT block. We also demonstrated that the extended AFTNet can also be applied to 1D data in Section 5.3. In addition, previous works define the loss in the magnitude image, while we calculate the loss in the complex-valued image domain, which preserves the relations between the real and imaginary parts. The phase is then derived from the output of AFTNet, which is essential for several phase-based applications, such as flow quantification and fat–water separation.
@@
CVNNs, especially complex-valued convolutional networks [22, 21], have been studied for MRI reconstruction, but have mainly focused on simple tasks or only applied them to the image domain. We investigate the different impacts of complex-valued convolutional networks on the k-space and image domain and extend the application to accelerated reconstruction and denoised reconstruction, which are more clinically important. We also incorporate domain-manifold learning by adding domain transform blocks that determine the mapping between the k-space and the image domain instead of the conventional discrete Fourier transform. It is more robust to noise and signal non-ideality due to imperfect acquisition. We also extend the application of complex-valued convolutional networks to 1D MRS denoised reconstruction, which has not been studied in previous work. Our results indicate that AFTNet is able to effectively decrease the contribution of the noise in the FIDs while preserving the quantifiable signals in the reconstructed spectra.
@@
One remaining methodological limitation is that the FC layers used by AFTNet narrow the application to datasets with various image matrix sizes. Although the convolutional layers are not sensitive to the image matrix sizes and cropping/padding can be applied to match the desired sizes, the features of FC layers need to be selected carefully which requires further investigation. Another parameter that needs to be taken into account is the coil number. In this study, we selected especially 4-channel MRI data for convenience of data preprocessing, while deep learning-based coil combinations could be incorporated into the framework in future work. Furthermore, diffusion models are shown to be a powerful tool for image reconstruction across body regions and coil numbers [13]. However, the score-MRI we compared in this study does not demonstrate superior performance compared with AFTNet and the inference stage time is long. This is potential because the backbone of the score-MRI is still a real-valued UNet structure and the relation between the real and imaginary part is not considered during the calculation of the score function. For future works, the AFTNet could be further extended by leveraging diffusion-based models with complex-valued convolutional networks as the backbone and careful optimization to reduce the inference time.